OpenStudy (babyx3boo):

at a barbeque there were 250 dinners served. children's plate were $1.50 each and adults plates were 2.00 each. if the total money collected was $441, how many of each type of plate was served?

zepdrix (zepdrix):

Hey :) Let \(\large\rm C\) represent `the number of child plates`. Then the `money collected from all of the children's plates` would be \(\large\rm 1.50C\). The number of plates C times 1.50 per plate. Also, let \(\large\rm A\) represent `the number of adult plates`. Then the `money collected from all of the adult plates` would be \(\large\rm 2.00A\). The number of plates A times 2.00 per plate.

zepdrix (zepdrix):

The total `number of plates` sold was 250. Adult plates plus child plates equals our total. \[\large\rm A+C=250\]

zepdrix (zepdrix):

The total `money collected` was 441. Adult money collected plus children money collected equals our total.\[\large\rm 2A+1.5C=441\]

zepdrix (zepdrix):

This gives us a system of `two equations` with `two unknowns`

zepdrix (zepdrix):

You can solve the system using a variety of methods. You're probably familiar with either `substitution` or `elimination`.

zepdrix (zepdrix):

Confused? :o

OpenStudy (babyx3boo):

A little, so there are two equations we need to figure out?

zepdrix (zepdrix):

There are two variables you need to figure out, A and C. You have to combine the equations in same way.

OpenStudy (babyx3boo):

that's where i need help

zepdrix (zepdrix):

\[\large\rm A+C=250\]\[\large\rm 2A+1.5C=441\] Substitution or Elimnation? :d No preference?

OpenStudy (babyx3boo):

oh wait i think i know now!

OpenStudy (babyx3boo):

Let's do elimination

zepdrix (zepdrix):

To do elimination, we need one of the variables to `match up` in both equations. It's much easier to match to the larger number, rather than trying to go backwards. So, notice that the second equation has a 2A. Our first equation has only A. We need to multiply the first equation by something to match up the A's.

OpenStudy (babyx3boo):

yess i did that she leads us to 2A+2C=4.00

OpenStudy (babyx3boo):

then i change the signs of the bottom ones

zepdrix (zepdrix):

Woops! :O That looks a little off.\[\large\rm 2(A+C=250)\]\[\large\rm 2A+2C=500\]Something like that, ya?

OpenStudy (babyx3boo):

yeah that's what i meant is that i was looking at the 2.00 of adults sorry

zepdrix (zepdrix):

\[\large\rm \quad~2A+2C~~~=500\]\[\large\rm -(2A+1.5C=441)\]----------------------- And you wanted to negate the second equation like this? And then combine, ya? Make sure you change the sign of EVERYTHING in the second equation, not just the A's :)

OpenStudy (babyx3boo):

Yess and then i got C= 118?

zepdrix (zepdrix):

Good! :) Any ideas on how to find A?

OpenStudy (babyx3boo):

yesss and i got A= 132

zepdrix (zepdrix):

Yayyy good job! ୧ʕ•̀ᴥ•́ʔ୨

zepdrix (zepdrix):

132 adult plates were served, 118 children plates were served

OpenStudy (babyx3boo):

Thank you :)